I dont get how to solve (x^2)-x-5=-5
I am stuck where (x^2)-x=0
Thanks for your help.

Answer by ptaylor(2198)

OK. We are down to:
x^2-x=0 Factor out an x from the left side:
x(x-1)=0
Now we have:
x=0
and
x=1
CK:
x^2-x-5=-5
x=0
0-0-5=-5
-5=-5-----OK
x=1
1-1-5=-5
-5=-5---ok
Hope this helps---ptaylor